\begin{thebibliography}{99}

%\cite{Smeenk:2018dbt}
\bibitem{Smeenk:2018dbt}
C.~Smeenk,
``Inflation and the origins of structure,''
Einstein Stud.\  \textbf{14}, 205-241 (2018)
%doi:10.1007/978-1-4939-7708-6_9

%\cite{Brout:1977ix}
\bibitem{Brout:1977ix}
R.~Brout, F.~Englert and E.~Gunzig,
``The creation of the universe as a quantum phenomenon,''
Annals Phys.\  \textbf{115}, 78 (1978)
%doi:10.1016/0003-4916(78)90176-8

%\cite{Starobinsky:1980te}
\bibitem{Starobinsky:1980te} 
  A.~A.~Starobinsky,
  ``A new type of isotropic cosmological models without singularity,''
  Phys.\ Lett.\  {\bf 91B}, 99 (1980)
  [Adv.\ Ser.\ Astrophys.\ Cosmol.\  {\bf 3}, 130 (1987)].
  %doi:10.1016/0370-2693(80)90670-X

%\cite{Sato:1980yn}
\bibitem{Sato:1980yn} 
  K.~Sato,
  ``First order phase transition of a vacuum and expansion of the universe,''
  Mon.\ Not.\ Roy.\ Astron.\ Soc.\  {\bf 195}, 467 (1981).

%\cite{Fang:1980wi}
\bibitem{Fang:1980wi} 
  L.~Z.~Fang,
  ``Entropy generation in the early universe by dissipative processes near the Higgs' phase transitions,''
  Phys.\ Lett.\  {\bf 95B}, 154 (1980).
  %doi:10.1016/0370-2693(80)90421-9

%\cite{Hawking:1969sw}
\bibitem{Hawking:1969sw} 
  S.~W.~Hawking and R.~Penrose,
  ``The singularities of gravitational collapse and cosmology,''
  Proc.\ Roy.\ Soc.\ Lond.\ A {\bf 314}, 529 (1970).
  %doi:10.1098/rspa.1970.0021

%\cite{Finelli:2001sr}
\bibitem{Finelli:2001sr} 
  F.~Finelli and R.~Brandenberger,
  ``On the generation of a scale invariant spectrum of adiabatic fluctuations in cosmological models with a contracting phase,''
  Phys.\ Rev.\ D {\bf 65}, 103522 (2002)
  %doi:10.1103/PhysRevD.65.103522
  [hep-th/0112249].

%\cite{Gasperini:1992em}
\bibitem{Gasperini:1992em} 
  M.~Gasperini and G.~Veneziano,
  ``Pre - big bang in string cosmology,''
  Astropart.\ Phys.\  {\bf 1}, 317 (1993)
  %doi:10.1016/0927-6505(93)90017-8
  [hep-th/9211021].

%\cite{Khoury:2001wf}
\bibitem{Khoury:2001wf} 
  J.~Khoury, B.~A.~Ovrut, P.~J.~Steinhardt and N.~Turok,
  ``The ekpyrotic universe: Colliding branes and the origin of the hot big bang,''
  Phys.\ Rev.\ D {\bf 64}, 123522 (2001)
  %doi:10.1103/PhysRevD.64.123522
  [hep-th/0103239].

%\cite{Brandenberger:1988aj}
\bibitem{Brandenberger:1988aj} 
  R.~H.~Brandenberger and C.~Vafa,
  ``Superstrings in the early universe,''
  Nucl.\ Phys.\ B {\bf 316}, 391 (1989).
  %doi:10.1016/0550-3213(89)90037-0

%\cite{Steinhardt:2001st}
\bibitem{Steinhardt:2001st} 
  P.~J.~Steinhardt and N.~Turok,
  ``Cosmic evolution in a cyclic universe,''
  Phys.\ Rev.\ D {\bf 65}, 126003 (2002)
  %doi:10.1103/PhysRevD.65.126003
  [hep-th/0111098].

%\cite{Feng:2018qnx}
\bibitem{Feng:2018qnx} 
  X.~H.~Feng, H.~Huang, S.~L.~Li, H.~L\"u and H.~Wei,
  ``Cosmological time crystals from Einstein-cubic gravities,''
  arXiv:1807.01720 [hep-th].

%\cite{Bondi:1948qk}
\bibitem{Bondi:1948qk} 
  H.~Bondi and T.~Gold,
  ``The steady-state theory of the expanding universe,''
  Mon.\ Not.\ Roy.\ Astron.\ Soc.\  {\bf 108}, 252 (1948).
 
%\cite{Hoyle:1948zz}
\bibitem{Hoyle:1948zz} 
  F.~Hoyle,
  ``A new model for the expanding universe,''
  Mon.\ Not.\ Roy.\ Astron.\ Soc.\  {\bf 108}, 372 (1948).

%\cite{Ellis:2002we}
\bibitem{Ellis:2002we} 
  G.~F.~R.~Ellis and R.~Maartens,
  ``The emergent universe: Inflationary cosmology with no singularity,''
  Class.\ Quant.\ Grav.\  {\bf 21}, 223 (2004)
  %doi:10.1088/0264-9381/21/1/015
  [gr-qc/0211082].

%\cite{Ellis:2003qz}
\bibitem{Ellis:2003qz}
G.~F.~R.~Ellis, J.~Murugan and C.~G.~Tsagas,
``The emergent universe: An explicit construction,''
Class.\ Quant.\ Grav.\  {\bf 21}, no. 1, 233 (2004)
%doi:10.1088/0264-9381/21/1/016
[gr-qc/0307112].

%\cite{Gibbons:1987jt}
\bibitem{Gibbons:1987jt}
G.~W.~Gibbons,
``The entropy and stability of the universe,''
Nucl.\ Phys.\ B {\bf 292}, 784 (1987).
%doi:10.1016/0550-3213(87)90670-5

%\cite{Gibbons:1988bm}
\bibitem{Gibbons:1988bm}
G.~W.~Gibbons,
``Sobolev's inequality, Jensen's theorem and the mass and entropy of the universe,''
Nucl.\ Phys.\ B {\bf 310}, 636 (1988).
%doi:10.1016/0550-3213(88)90096-X

%\cite{Mulryne:2005ef}
\bibitem{Mulryne:2005ef} 
D.~J.~Mulryne, R.~Tavakol, J.~E.~Lidsey and G.~F.~R.~Ellis,
``An emergent universe from a loop,''
Phys.\ Rev.\ D {\bf 71}, 123512 (2005)
%doi:10.1103/PhysRevD.71.123512
[astro-ph/0502589].

%\cite{Eddington:1930zz}
\bibitem{Eddington:1930zz}
A.~S.~Eddington,
``On the instability of Einstein's spherical world,''
Mon.\ Not.\ Roy.\ Astron.\ Soc.\  {\bf 90}, 668 (1930).

%\cite{Barrow:2003ni}
\bibitem{Barrow:2003ni}
J.~D.~Barrow, G.~F.~R.~Ellis, R.~Maartens and C.~G.~Tsagas,
``On the stability of the Einstein static universe,''
Class.\ Quant.\ Grav.\  {\bf 20}, L155 (2003)
%doi:10.1088/0264-9381/20/11/102
[gr-qc/0302094].

%\cite{delCampo:2007mp}
\bibitem{delCampo:2007mp}
S.~del Campo, R.~Herrera and P.~Labrana,
``Emergent universe in a Jordan-Brans-Dicke theory,''
JCAP {\bf 0711}, 030 (2007)
%doi:10.1088/1475-7516/2007/11/030
[arXiv:0711.1559 [gr-qc]].

%\cite{delCampo:2009kp}
\bibitem{delCampo:2009kp}
S.~del Campo, R.~Herrera and P.~Labrana,
``On the stability of Jordan-Brans-Dicke static universe,''
JCAP {\bf 0907}, 006 (2009)
%doi:10.1088/1475-7516/2009/07/006
[arXiv:0905.0614 [gr-qc]].

%\cite{Huang:2014fia}
\bibitem{Huang:2014fia}
H.~Huang, P.~Wu and H.~Yu,
``Stability of the Einstein static universe in the Jordan-Brans-Dicke theory,''
Phys.\ Rev.\ D {\bf 89}, no. 10, 103521 (2014).
%doi:10.1103/PhysRevD.89.103521

%\cite{Huang:2018kqr}
\bibitem{Huang:2018kqr} 
  Q.~Huang, P.~Wu and H.~Yu,
  ``Stability of Einstein static universe in gravity theory with a non-minimal derivative coupling,''
  Eur.\ Phys.\ J.\ C {\bf 78}, no. 1, 51 (2018)
  %doi:10.1140/epjc/s10052-018-5533-6
  [arXiv:1801.04358 [gr-qc]].
  
%\cite{Miao:2016obc}
\bibitem{Miao:2016obc}
  H.~Miao, P.~Wu and H.~Yu,
  ``Stability of the Einstein static Universe in the scalar–tensor theory of gravity,''
  Class.\ Quant.\ Grav.\  {\bf 33}, no. 21, 215011 (2016).
%  doi:10.1088/0264-9381/33/21/215011

  %\cite{Atazadeh:2015zma}
\bibitem{Atazadeh:2015zma}
  K.~Atazadeh and F.~Darabi,
  ``Einstein static Universe in non-minimal kinetic coupled gravity,''
  Phys.\ Lett.\ B {\bf 744}, 363 (2015)
 % doi:10.1016/j.physletb.2015.04.022
  [arXiv:1502.06721 [gr-qc]].

%\cite{Parisi:2012cg}
\bibitem{Parisi:2012cg}
L.~Parisi, N.~Radicella and G.~Vilasi,
``On the stability of the Einstein Static Universe in Massive Gravity,''
Phys.\ Rev.\ D {\bf 86}, 024035 (2012)
%doi:10.1103/PhysRevD.86.024035
[arXiv:1207.3922 [gr-qc]].

%\cite{Li:2019laq}
\bibitem{Li:2019laq} 
  S.~L.~Li, H.~Lü, H.~Wei, P.~Wu and H.~Yu,
  ``Emergent universe scenario, bouncing universes, and cyclic universes in degenerate massive gravity,''
  Phys.\ Rev.\ D {\bf 99}, no. 10, 104057 (2019)
  %doi:10.1103/PhysRevD.99.104057
  [arXiv:1903.03940 [gr-qc]].

%\cite{Mousavi:2016eof}
\bibitem{Mousavi:2016eof} 
  M.~Mousavi and F.~Darabi,
  ``On the stability of Einstein static universe at background level in massive bigravity,''
  Nucl.\ Phys.\ B {\bf 919}, 523 (2017)
  %doi:10.1016/j.nuclphysb.2017.04.002
  [arXiv:1607.04377 [gr-qc]].


%\cite{Boehmer:2007tr}
\bibitem{Boehmer:2007tr}
C.~G.~Boehmer, L.~Hollenstein and F.~S.~N.~Lobo,
``Stability of the Einstein static universe in f(R) gravity,''
Phys.\ Rev.\ D {\bf 76}, 084005 (2007)
%doi:10.1103/PhysRevD.76.084005
[arXiv:0706.1663 [gr-qc]].

%\cite{Goheer:2008tn}
\bibitem{Goheer:2008tn}
N.~Goheer, R.~Goswami and P.~K.~S.~Dunsby,
``Dynamics of f(R)-cosmologies containing Einstein static models,''
Class.\ Quant.\ Grav.\  {\bf 26}, 105003 (2009)
doi:10.1088/0264-9381/26/10/105003
[arXiv:0809.5247 [gr-qc]].
%%CITATION = doi:10.1088/0264-9381/26/10/105003;%%
%44 citations counted in INSPIRE as of 17 Apr 2017

%\cite{Seahra:2009ft}
\bibitem{Seahra:2009ft}
S.~S.~Seahra and C.~G.~Boehmer,
``Einstein static universes are unstable in generic f(R) models,''
Phys.\ Rev.\ D {\bf 79}, 064009 (2009)
doi:10.1103/PhysRevD.79.064009
[arXiv:0901.0892 [gr-qc]].
%%CITATION = doi:10.1103/PhysRevD.79.064009;%%
%41 citations counted in INSPIRE as of 17 Apr 2017

%\cite{Wu:2011xa}
\bibitem{Wu:2011xa}
P.~Wu and H.~Yu,
``The Stability of the Einstein static state in $f(T)$ gravity,''
Phys.\ Lett.\ B {\bf 703}, 223 (2011)
%doi:10.1016/j.physletb.2011.07.087
[arXiv:1108.5908 [gr-qc]].

%\cite{Li:2013xea}
\bibitem{Li:2013xea}
J.~T.~Li, C.~C.~Lee and C.~Q.~Geng,
``Einstein Static Universe in Exponential $f(T)$ Gravity,''
Eur.\ Phys.\ J.\ C {\bf 73}, no. 2, 2315 (2013)
%doi:10.1140/epjc/s10052-013-2315-z
[arXiv:1302.2688 [gr-qc]].

%\cite{Bohmer:2009fc}
\bibitem{Bohmer:2009fc}
C.~G.~Boehmer and F.~S.~N.~Lobo,
``Stability of the Einstein static universe in modified Gauss-Bonnet gravity,''
Phys.\ Rev.\ D {\bf 79}, 067504 (2009)
%doi:10.1103/PhysRevD.79.067504
[arXiv:0902.2982 [gr-qc]].

%\cite{Huang:2015kca}
\bibitem{Huang:2015kca}
H.~Huang, P.~Wu and H.~Yu,
``Instability of the Einstein static universe in modified Gauss-Bonnet gravity,''
Phys.\ Rev.\ D {\bf 91}, no. 2, 023507 (2015).
%doi:10.1103/PhysRevD.91.023507

%\cite{Paul:2008id}
\bibitem{Paul:2008id} 
  B.~C.~Paul and S.~Ghose,
  ``Emergent Universe Scenario in the Einstein-Gauss-Bonnet Gravity with Dilaton,''
  Gen.\ Rel.\ Grav.\  {\bf 42}, 795 (2010)
  %doi:10.1007/s10714-009-0880-0
  [arXiv:0809.4131 [hep-th]].


%\cite{Wu:2009ah}
\bibitem{Wu:2009ah}
P.~Wu and H.~W.~Yu,
``Emergent universe from the Horava-Lifshitz gravity,''
Phys.\ Rev.\ D {\bf 81}, 103522 (2010)
doi:10.1103/PhysRevD.81.103522
[arXiv:0909.2821 [gr-qc]].
%%CITATION = doi:10.1103/PhysRevD.81.103522;%%
%70 citations counted in INSPIRE as of 17 Apr 2017

%\cite{Boehmer:2009yz}
\bibitem{Boehmer:2009yz}
C.~G.~Boehmer and F.~S.~N.~Lobo,
``Stability of the Einstein static universe in IR modified Horava gravity,''
Eur.\ Phys.\ J.\ C {\bf 70}, 1111 (2010)
doi:10.1140/epjc/s10052-010-1503-3
[arXiv:0909.3986 [gr-qc]].
%%CITATION = doi:10.1140/epjc/s10052-010-1503-3;%%
%74 citations counted in INSPIRE as of 17 Apr 2017

%\cite{Heydarzade:2015hra}
\bibitem{Heydarzade:2015hra}
Y.~Heydarzade, M.~Khodadi and F.~Darabi,
``Deformed Hořava–Lifshitz cosmology and stability of the Einstein static universe,''
Theor.\ Math.\ Phys.\  {\bf 190}, no. 1, 130 (2017)
[Teor.\ Mat.\ Fiz.\  {\bf 190}, no. 1, 150 (2017)]
%doi:10.4213/tmf9066, 10.1134/S0040577917010111
[arXiv:1502.04445 [gr-qc]].

%\cite{Goswami:2008fs}
\bibitem{Goswami:2008fs}
  R.~Goswami, N.~Goheer and P.~K.~S.~Dunsby,
  ``The Existence of Einstein Static Universes and their Stability in Fourth order Theories of Gravity,''
  Phys.\ Rev.\ D {\bf 78}, 044011 (2008)
 % doi:10.1103/PhysRevD.78.044011
  [arXiv:0804.3528 [gr-qc]].

%\cite{Gergely:2001tn}
\bibitem{Gergely:2001tn}
  L.~A.~Gergely and R.~Maartens,
  ``Brane world generalizations of the Einstein static universe,''
  Class.\ Quant.\ Grav.\  {\bf 19}, 213 (2002)
 % doi:10.1088/0264-9381/19/2/303
  [gr-qc/0105058].

%\cite{Zhang:2010qwa}
\bibitem{Zhang:2010qwa}
  K.~Zhang, P.~Wu and H.~W.~Yu,
  ``The Stability of Einstein static universe in the DGP braneworld,''
  Phys.\ Lett.\ B {\bf 690}, 229 (2010)
  %doi:10.1016/j.physletb.2010.05.040
  [arXiv:1005.4201 [gr-qc]].

  %\cite{Atazadeh:2014xsa}
\bibitem{Atazadeh:2014xsa}
  K.~Atazadeh, Y.~Heydarzade and F.~Darabi,
  ``Einstein Static Universe in Braneworld Scenario,''
  Phys.\ Lett.\ B {\bf 732}, 223 (2014)
 % doi:10.1016/j.physletb.2014.03.009
  [arXiv:1401.7638 [gr-qc]].

  %\cite{Zhang:2016obw}
\bibitem{Zhang:2016obw}
  K.~Zhang, P.~Wu, H.~Yu and L.~W.~Luo,
  ``Stability of Einstein static state universe in the spatially flat branemodels,''
  Phys.\ Lett.\ B {\bf 758}, 37 (2016).
%  doi:10.1016/j.physletb.2016.04.049

   
  %\cite{Boehmer:2013oxa}
\bibitem{Boehmer:2013oxa}
  C.~G.~Boehmer, F.~S.~N.~Lobo and N.~Tamanini,
  ``Einstein static Universe in hybrid metric-Palatini gravity,''
  Phys.\ Rev.\ D {\bf 88}, 104019 (2013)
%  doi:10.1103/PhysRevD.88.104019
  [arXiv:1305.0025 [gr-qc]].

    %\cite{Boehmer:2003iv}
\bibitem{Boehmer:2003iv}
  C.~G.~Boehmer,
 ``The Einstein static universe with torsion and the sign problem of the cosmological constant,''
  Class.\ Quant.\ Grav.\  {\bf 21}, 1119 (2004)
  %doi:10.1088/0264-9381/21/4/025
  [gr-qc/0310058].
 

%\cite{Canonico:2010fd}
\bibitem{Canonico:2010fd}
  R.~Canonico and L.~Parisi,
  ``Stability of the Einstein Static Universe in open cosmological models,''
  Phys.\ Rev.\ D {\bf 82}, 064005 (2010)
 % doi:10.1103/PhysRevD.82.064005
  [arXiv:1005.3673 [gr-qc]].

%\cite{Huang:2015zma}
\bibitem{Huang:2015zma} 
  Q.~Huang, P.~Wu and H.~Yu,
  ``Emergent scenario in the Einstein-Cartan theory,''
  Phys.\ Rev.\ D {\bf 91}, no. 10, 103502 (2015)
  %doi:10.1103/PhysRevD.91.103502
  [arXiv:1504.05284 [gr-qc]].

  %\cite{Carneiro:2009et}
\bibitem{Carneiro:2009et}
  S.~Carneiro and R.~Tavakol,
  ``Stability of the Einstein static universe in presence of vacuum energy,''
  Phys.\ Rev.\ D {\bf 80}, 043528 (2009)
  %doi:10.1103/PhysRevD.80.043528
  [arXiv:0907.4795 [astro-ph.CO]].

%\cite{Atazadeh:2016yeh}
\bibitem{Atazadeh:2016yeh} 
  K.~Atazadeh and F.~Darabi,
  ``Einstein static universe from GUP,''
  Phys.\ Dark Univ.\  {\bf 16}, 87 (2017)
 % doi:10.1016/j.dark.2017.04.008
  [arXiv:1701.00060 [gr-qc]].
 
 
%\cite{Li:2017ttl}
\bibitem{Li:2017ttl} 
  S.~L.~Li and H.~Wei,
  ``Stability of the Einstein static universe in Eddington-inspired Born-Infeld theory,''
  Phys.\ Rev.\ D {\bf 96}, no. 2, 023531 (2017)
 % doi:10.1103/PhysRevD.96.023531
  [arXiv:1705.06819 [gr-qc]].

%\cite{Parisi:2007kv}
\bibitem{Parisi:2007kv}
L.~Parisi, M.~Bruni, R.~Maartens and K.~Vandersloot,
``The Einstein static universe in Loop Quantum Cosmology,''
Class.\ Quant.\ Grav.\  {\bf 24}, 6243 (2007)
%doi:10.1088/0264-9381/24/24/007
[arXiv:0706.4431 [gr-qc]].

%\cite{Yu:2019cku}
\bibitem{Yu:2019cku} 
  Z.~X.~Yu and H.~Wei,
  ``Lemaitre-Tolman-Bondi Static Universe in Rastall-like Gravity,''
  arXiv:1907.12517 [gr-qc].

%\cite{Zhang:2013ykz}
\bibitem{Zhang:2013ykz} 
  K.~Zhang, P.~Wu and H.~Yu,
  ``Emergent universe in spatially flat cosmological model,''
  JCAP {\bf 1401}, 048 (2014)
  %doi:10.1088/1475-7516/2014/01/048
  [arXiv:1311.4051 [gr-qc]].

%\cite{Lovelock:1971yv}
\bibitem{Lovelock:1971yv} 
  D.~Lovelock,
  ``The Einstein tensor and its generalizations,''
  J.\ Math.\ Phys.\  {\bf 12}, 498 (1971).
 % doi:10.1063/1.1665613

%\cite{Boulware:1985wk}
\bibitem{Boulware:1985wk}
D.~G.~Boulware and S.~Deser,
``String generated gravity models,''
Phys.\ Rev.\ Lett.\  \textbf{55}, 2656 (1985)
%doi:10.1103/PhysRevLett.55.2656

%\cite{Cai:2001dz}
\bibitem{Cai:2001dz}
R.~Cai,
``Gauss-Bonnet black holes in AdS spaces,''
Phys.\ Rev.\ D \textbf{65}, 084014 (2002)
%doi:10.1103/PhysRevD.65.084014
[arXiv:hep-th/0109133 [hep-th]].

%\cite{Cvetic:2001bk}
\bibitem{Cvetic:2001bk} 
  M.~Cvetic, S.~Nojiri and S.~D.~Odintsov,
  ``Black hole thermodynamics and negative entropy in de Sitter and anti-de Sitter Einstein-Gauss-Bonnet gravity,''
  Nucl.\ Phys.\ B {\bf 628}, 295 (2002)
  %doi:10.1016/S0550-3213(02)00075-5
  [hep-th/0112045].

%\cite{Brigante:2007nu}
\bibitem{Brigante:2007nu} 
  M.~Brigante, H.~Liu, R.~C.~Myers, S.~Shenker and S.~Yaida,
  ``Viscosity bound violation in higher derivative gravity,''
  Phys.\ Rev.\ D {\bf 77}, 126006 (2008)
  %doi:10.1103/PhysRevD.77.126006
  [arXiv:0712.0805 [hep-th]].

%\cite{Pan:2009xa}
\bibitem{Pan:2009xa} 
  Q.~Pan, B.~Wang, E.~Papantonopoulos, J.~Oliveira and A.~B.~Pavan,
  ``Holographic superconductors with various condensates in Einstein-Gauss-Bonnet gravity,''
  Phys.\ Rev.\ D {\bf 81}, 106007 (2010)
  %doi:10.1103/PhysRevD.81.106007
  [arXiv:0912.2475 [hep-th]].

%\cite{Charmousis:2002rc}
\bibitem{Charmousis:2002rc} 
  C.~Charmousis and J.~F.~Dufaux,
  ``General Gauss-Bonnet brane cosmology,''
  Class.\ Quant.\ Grav.\  {\bf 19}, 4671 (2002)
  %doi:10.1088/0264-9381/19/18/304
  [hep-th/0202107].

%\cite{Nojiri:2005jg}
\bibitem{Nojiri:2005jg} 
  S.~Nojiri and S.~D.~Odintsov,
  ``Modified Gauss-Bonnet theory as gravitational alternative for dark energy,''
  Phys.\ Lett.\ B {\bf 631}, 1 (2005)
  %doi:10.1016/j.physletb.2005.10.010
  [hep-th/0508049].

%\cite{Nojiri:2005vv}
\bibitem{Nojiri:2005vv} 
  S.~Nojiri, S.~D.~Odintsov and M.~Sasaki,
  ``Gauss-Bonnet dark energy,''
  Phys.\ Rev.\ D {\bf 71}, 123509 (2005)
  %doi:10.1103/PhysRevD.71.123509
  [hep-th/0504052].
 
%\cite{Glavan:2019inb}
\bibitem{Glavan:2019inb} 
  D.~Glavan and C.~Lin,
  ``Einstein-Gauss-Bonnet gravity in 4-dimensional space-time,''
  Phys.\ Rev.\ Lett.\  {\bf 124}, no. 8, 081301 (2020)
  %doi:10.1103/PhysRevLett.124.081301
  [arXiv:1905.03601 [gr-qc]].

%\cite{Tomozawa:2011gp}
\bibitem{Tomozawa:2011gp} 
  Y.~Tomozawa,
  ``Quantum corrections to gravity,''
  arXiv:1107.1424 [gr-qc].
  
%\cite{Cognola:2013fva}
\bibitem{Cognola:2013fva}
G.~Cognola, R.~Myrzakulov, L.~Sebastiani and S.~Zerbini,
``Einstein gravity with Gauss-Bonnet entropic corrections,''
Phys.\ Rev.\ D \textbf{88}, no.2, 024006 (2013)
%doi:10.1103/PhysRevD.88.024006
[arXiv:1304.1878 [gr-qc]].

%\cite{Konoplya:2020bxa}
\bibitem{Konoplya:2020bxa} 
  R.~A.~Konoplya and A.~F.~Zinhailo,
  ``Quasinormal modes, stability and shadows of a black hole in the novel 4D Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.01188 [gr-qc].

%\cite{Guo:2020zmf}
\bibitem{Guo:2020zmf} 
  M.~Guo and P.~C.~Li,
  ``The innermost stable circular orbit and shadow in the novel $4D$ Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.02523 [gr-qc].

%\cite{Fernandes:2020rpa}
\bibitem{Fernandes:2020rpa} 
  P.~G.~S.~Fernandes,
  ``Charged Black Holes in AdS Spaces in $4D$ Einstein Gauss-Bonnet Gravity,''
  arXiv:2003.05491 [gr-qc].

%\cite{Casalino:2020kbt}
\bibitem{Casalino:2020kbt} 
  A.~Casalino, A.~Colleaux, M.~Rinaldi and S.~Vicentini,
  ``Regularized Lovelock gravity,''
  arXiv:2003.07068 [gr-qc].

\bibitem{Wei:2020ght} 
  S.~W.~Wei and Y.~X.~Liu,
  ``Testing the nature of Gauss-Bonnet gravity by four-dimensional rotating black hole shadow,''
  arXiv:2003.07769 [gr-qc].

%\cite{Konoplya:2020qqh}
\bibitem{Konoplya:2020qqh} 
  R.~A.~Konoplya and A.~Zhidenko,
  ``Black holes in the four-dimensional Einstein-Lovelock gravity,''
  arXiv:2003.07788 [gr-qc].

%\cite{Kumar:2020owy}
\bibitem{Kumar:2020owy} 
  R.~Kumar and S.~G.~Ghosh,
  ``Rotating black holes in the novel $4D$ Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.08927 [gr-qc].

%\cite{Hegde:2020xlv}
\bibitem{Hegde:2020xlv} 
  K.~Hegde, A.~Naveena Kumara, C.~L.~A.~Rizwan, A.~K.~M. and M.~S.~Ali,
  ``Thermodynamics, Phase Transition and Joule Thomson Expansion of novel 4-D Gauss Bonnet AdS Black Hole,''
  arXiv:2003.08778 [gr-qc].

%\cite{Doneva:2020ped}
\bibitem{Doneva:2020ped} 
  D.~D.~Doneva and S.~S.~Yazadjiev,
  ``Relativistic stars in 4D Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.10284 [gr-qc].
 
%\cite{Ghosh:2020vpc}
\bibitem{Ghosh:2020vpc} 
  S.~G.~Ghosh and S.~D.~Maharaj,
  ``Radiating black holes in the novel 4D Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.09841 [gr-qc].

%\cite{Zhang:2020qew}
\bibitem{Zhang:2020qew} 
  Y.~P.~Zhang, S.~W.~Wei and Y.~X.~Liu,
  ``Spinning test particle in four-dimensional Einstein-Gauss-Bonnet Black Hole,''
  arXiv:2003.10960 [gr-qc].

%\cite{Lu:2020iav}
\bibitem{Lu:2020iav} 
  H.~Lu and Y.~Pang,
  ``Horndeski Gravity as $D\rightarrow4$ Limit of Gauss-Bonnet,''
  arXiv:2003.11552 [gr-qc].

%\cite{Singh:2020xju}
\bibitem{Singh:2020xju} 
  D.~V.~Singh and S.~Siwach,
  ``Thermodynamics and P-v criticality of Bardeen-AdS Black Hole in 4-D Einstein-Gauss-Bonnet Gravity,''
  arXiv:2003.11754 [gr-qc].

%\cite{Konoplya:2020ibi}
\bibitem{Konoplya:2020ibi} 
  R.~A.~Konoplya and A.~Zhidenko,
  ``BTZ black holes with higher curvature corrections in the 3D Einstein-Lovelock theory,''
  arXiv:2003.12171 [gr-qc].

%\cite{Ghosh:2020syx}
\bibitem{Ghosh:2020syx} 
  S.~G.~Ghosh and R.~Kumar,
  ``Generating black holes in the novel $4D$ Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.12291 [gr-qc].

%\cite{Konoplya:2020juj}
\bibitem{Konoplya:2020juj} 
  R.~A.~Konoplya and A.~Zhidenko,
  ``(In)stability of black holes in the 4D Einstein-Gauss-Bonnet and Einstein-Lovelock gravities,''
  arXiv:2003.12492 [gr-qc].

%\cite{Kobayashi:2020wqy}
\bibitem{Kobayashi:2020wqy} 
  T.~Kobayashi,
  ``Effective scalar-tensor description of regularized Lovelock gravity in four dimensions,''
  arXiv:2003.12771 [gr-qc].

%\cite{Zhang:2020qam}
\bibitem{Zhang:2020qam} 
  C.~Y.~Zhang, P.~C.~Li and M.~Guo,
  ``Greybody factor and power spectra of the Hawking radiation in the novel $4D$ Einstein-Gauss-Bonnet de-Sitter gravity,''
  arXiv:2003.13068 [hep-th].

%\cite{HosseiniMansoori:2020yfj}
\bibitem{HosseiniMansoori:2020yfj} 
  S.~A.~Hosseini Mansoori,
  ``Thermodynamic geometry of novel 4-D Gauss Bonnet AdS Black Hole,''
  arXiv:2003.13382 [gr-qc].

%\cite{Kumar:2020uyz}
\bibitem{Kumar:2020uyz} 
  A.~Kumar and R.~Kumar,
  ``Bardeen black holes in the novel $4D$ Einstein-Gauss-Bonnet gravity,''
  arXiv:2003.13104 [gr-qc].

%\cite{Singh:2020nwo}
\bibitem{Singh:2020nwo} 
  D.~V.~Singh, S.~G.~Ghosh and S.~D.~Maharaj,
  ``Clouds of string in $4D$ novel Einstein-Gauss-Bonnet black holes,''
  arXiv:2003.14136 [gr-qc].

%\cite{Wei:2020poh}
\bibitem{Wei:2020poh} 
  S.~W.~Wei and Y.~X.~Liu,
  ``Extended thermodynamics and microstructures of four-dimensional charged Gauss-Bonnet black hole in AdS space,''
  arXiv:2003.14275 [gr-qc].

%\cite{Churilova:2020aca}
\bibitem{Churilova:2020aca} 
  M.~S.~Churilova,
  ``Quasinormal modes of the Dirac field in the novel 4D Einstein-Gauss-Bonnet gravity,''
  arXiv:2004.00513 [gr-qc].

%\cite{Kumar:2020xvu}
\bibitem{Kumar:2020xvu} 
  A.~Kumar and S.~G.~Ghosh,
  ``Hayward black holes in the novel $4D$ Einstein-Gauss-Bonnet gravity,''
  arXiv:2004.01131 [gr-qc].

%\cite{Islam:2020xmy}
\bibitem{Islam:2020xmy} 
  S.~U.~Islam, R.~Kumar and S.~G.~Ghosh,
  ``Gravitational lensing by black holes in $4D$ Einstein-Gauss-Bonnet gravity,''
  arXiv:2004.01038 [gr-qc].

%\cite{Liu:2020vkh}
\bibitem{Liu:2020vkh}
C.~Liu, T.~Zhu and Q.~Wu,
``Thin Accretion Disk around a four-dimensional Einstein-Gauss-Bonnet Black Hole,''
[arXiv:2004.01662 [gr-qc]].

%\cite{Heydari-Fard:2020sib}
\bibitem{Heydari-Fard:2020sib}
M.~Heydari-Fard, M.~Heydari-Fard and H.~Sepangi,
``Bending of light in novel 4$D$ Gauss-Bonnet-de Sitter black holes by Rindler-Ishak method,''
[arXiv:2004.02140 [gr-qc]].

%\cite{Konoplya:2020cbv}
\bibitem{Konoplya:2020cbv}
R.~A.~Konoplya and A.~F.~Zinhailo,
``Grey-body factors and Hawking radiation of black holes in $4D$ Einstein-Gauss-Bonnet gravity,''
[arXiv:2004.02248 [gr-qc]].

%\cite{Jin:2020emq}
\bibitem{Jin:2020emq}
X.~Jin, Y.~Gao and D.~Liu,
``Strong gravitational lensing of a 4D Einstein-Gauss-Bonnet black hole in homogeneous plasma,''
[arXiv:2004.02261 [gr-qc]].

%\cite{Ai:2020peo}
\bibitem{Ai:2020peo}
W.~Ai,
``A note on the novel 4D Einstein-Gauss-Bonnet gravity,''
[arXiv:2004.02858 [gr-qc]].

%\cite{Zhang:2020sjh}
\bibitem{Zhang:2020sjh}
C.~Zhang, S.~Zhang, P.~Li and M.~Guo,
``Superradiance and stability of the novel 4D charged Einstein-Gauss-Bonnet black hole,''
[arXiv:2004.03141 [gr-qc]].

%\cite{EslamPanah:2020hoj}
\bibitem{EslamPanah:2020hoj}
B.~Eslam Panah and K.~Jafarzade,
``4D Einstein-Gauss-Bonnet AdS Black Holes as Heat Engine,''
[arXiv:2004.04058 [hep-th]].

%\cite{NaveenaKumara:2020rmi}
\bibitem{NaveenaKumara:2020rmi}
A.~Naveena Kumara, C.~A.~Rizwan, K.~Hegde, M.~S.~Ali and A.~K.~M,
``Rotating 4D Gauss-Bonnet black hole as particle accelerator,''
[arXiv:2004.04521 [gr-qc]].

%\cite{Gurses:2020ofy}
\bibitem{Gurses:2020ofy}
M.~Gurses, T.~C.~Sisman and B.~Tekin,
``Is there a novel Einstein-Gauss-Bonnet theory in four dimensions?,''
[arXiv:2004.03390 [gr-qc]].

%\cite{Harrison:1967zza}
\bibitem{Harrison:1967zza} 
E.~R.~Harrison,
``Normal Modes of Vibrations of the Universe,''
Rev.\ Mod.\ Phys.\  {\bf 39}, 862 (1967).
%doi:10.1103/RevModPhys.39.862



\end{thebibliography}

\documentclass[11pt]{article}
\pdfoutput=1
\usepackage{amsmath,amssymb,color,epsfig,cite}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{setspace}
\usepackage{amsthm,mathenv}

\allowdisplaybreaks[4]

\def\baselinestretch{1.4}
\usepackage{amsfonts}

\makeatletter
\@addtoreset{equation}{section}
\makeatother

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bea}{\setlength\arraycolsep{2pt} \begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\nn}{\nonumber}

\def\0{{\sst{(0)}}}
\def\1{{\sst{(1)}}}
\def\2{{\sst{(2)}}}
\def\3{{\sst{(3)}}}
\def\4{{\sst{(4)}}}
\def\5{{\sst{(5)}}}
\def\6{{\sst{(6)}}}
\def\7{{\sst{(7)}}}
\def\8{{\sst{(8)}}}
\def\sst#1{{\scriptscriptstyle #1}}

\begin{document}

\begin{center}
{\large {\bf Stability of the Einstein Static Universe in $4 D$ Gauss-Bonnet Gravity}}

\vspace{15pt}
{\large Shou-Long Li,  Puxun Wu and Hongwei Yu}

\vspace{15pt}

{\it Department of Physics and Synergetic Innovation Center for Quantum Effect and Applications, Hunan Normal University, Changsha 410081, China}

\vspace{40pt}

\underline{ABSTRACT}
\end{center}
By rescaling the Gauss-Bonnet (GB) coupling constant $\alpha \rightarrow \alpha/(D-4)$ and considering the $D \rightarrow 4$ limit, the GB gravity gives rise to nontrivial modification of general relativity in four dimensions. In this work, we investigate the realization of the emergent universe scenario in the $4 D$ GB gravity. First, we obtain the Einstein static universe filled with a perfect fluid. Then, we show that  both spatially closed and open universes can be stable against both homogeneous and inhomogeneous scalar perturbations simultaneously.

\vfill
 shoulongli@hunnu.edu.cn \ \ \ pxwu@hunnu.edu.cn \ \ \ hwyu@hunnu.edu.cn


\thispagestyle{empty}

\pagebreak



\newpage

\section{Introduction}\label{sec1}
According to the standard big bang cosmology, our Universe began with a big bang and evolved from a nearly infinitely hot dense initial state. By postulating that the universe expands exponentially in the very early stage of its evolution, inflationary cosmology~\cite{Smeenk:2018dbt, Brout:1977ix, Starobinsky:1980te, Sato:1980yn, Fang:1980wi}  explains the horizon problem, flatness problem and magnetic-monopole problem very well. However, the singularity problem still remains one of the most prominent  open problems in the standard model of cosmology~\cite{Hawking:1969sw}. 
In order to avoid the initial singularity, many potential solutions have been proposed to provide a nonsingular cosmological model. One class of the solutions is a scenario which assumes that  the universe undergoes a transition from a contracting phase to an expanding phase~\cite{Finelli:2001sr, Gasperini:1992em, Khoury:2001wf, Brandenberger:1988aj, Steinhardt:2001st, Feng:2018qnx}. Another class  is the past-eternal universe which is a nonsingular cosmological model without a bounce~\cite{Bondi:1948qk, Hoyle:1948zz, Ellis:2002we, Ellis:2003qz}. One important scenario of the past-eternal universe is the so-called emergent universe  proposed by Ellis et al.~\cite{Ellis:2002we, Ellis:2003qz}, which assumes that  initially,  our Universe starts from an Einstein static universe with finite size, and then evolves to an inflationary era. This scenario has many attractive advantages. For instance, the horizon problem can be solved before inflation begins. In addition, there is no singularity, no exotic physics is involved, and the quantum gravity regime can even be avoided. Furthermore, it has been argued that the Einstein static state is favored by entropy considerations as the initial state for our universe~\cite{Gibbons:1987jt, Gibbons:1988bm, Mulryne:2005ef}.

The Einstein static universe was first proposed  by Einstein to describe a homogeneous and isotropic closed universe which neither expands nor contracts. However, it was soon demonstrated by Eddington~\cite{Eddington:1930zz} that the solution is unstable with respect to homogeneous and isotropic scalar perturbations in general relativity (GR).  On the other hand, it was also found that the universe is expanding rather than static. So, the mainstream cosmological model for the evolution of our universe  was replaced by a dynamical model described Friedmann-Robertson-Walker~(FRW) metric. Recently, The Einstein static universe was revived in the emergent universe scenario to avoid the big bang singularity. It should be pointed out that the  stability of the Einstein static universe is crucial to a successful realization of the scenario. Actually, it was found that the Einstein static universe is stable against small inhomogeneous vector and tensor perturbations as well as adiabatic scalar density perturbations if the universe contains a perfect fluid with $w = c_s^2 >1/5$ in GR~\cite{Gibbons:1987jt, Gibbons:1988bm, Barrow:2003ni}. The stability of the Einstein static universe has also been extensively studied in vast modified gravities, for example, GR with an extra field~\cite{delCampo:2007mp, delCampo:2009kp, Huang:2014fia,Huang:2018kqr, Miao:2016obc, Atazadeh:2015zma, Parisi:2012cg, Li:2019laq, Mousavi:2016eof}, higher derivative gravity~\cite{Boehmer:2007tr, Seahra:2009ft, Goheer:2008tn, Wu:2011xa, Li:2013xea, Paul:2008id, Bohmer:2009fc, Huang:2015kca, Wu:2009ah, Boehmer:2009yz, Heydarzade:2015hra, Goswami:2008fs}, high dimensional gravity~\cite{Zhang:2010qwa, Gergely:2001tn, Atazadeh:2014xsa, Zhang:2016obw} and so on~\cite{Boehmer:2013oxa, Huang:2015zma, Boehmer:2003iv, Canonico:2010fd, Carneiro:2009et,Atazadeh:2016yeh, Li:2017ttl, Parisi:2007kv, Yu:2019cku,Zhang:2013ykz}. 

Among these modified gravities, the Gauss-Bonnet (GB) gravity is the simplest modification of general relativity to include higher derivative curvature terms, while the resulting equation of motion (EOM) is still second order and has been proven to be free of ghosts~\cite{Lovelock:1971yv}. In past years, the GB gravity has been extensively studied in many aspects, such as  black holes and holography, e.g.~\cite{Boulware:1985wk, Cai:2001dz, Cvetic:2001bk, Brigante:2007nu, Pan:2009xa}. Since the GB combination is a topological invariant and plays no role in four dimensions,  the original GB gravity has been rarely studied in cosmology directly. Generally, the GB gravity was considered in high dimensional brane cosmology~\cite{Charmousis:2002rc}, and modified by the function $f({\cal G})$~\cite{Nojiri:2005jg}  or coupled with scalar fields~\cite{Nojiri:2005vv}. The emergent universe scenario has also been studied in the modified GB gravity~\cite{ Bohmer:2009fc, Huang:2015kca} and the GB gravity coupled with dilatons~\cite{Paul:2008id}. However, the Einstein static universe is unstable when both homogeneous and inhomogeneous scalar perturbations are taken into account although it is stable against only homogeneous or inhomogeneous perturbations. 

Recently, the GB gravity was revived in four dimensions by rescaling the GB coupling constant and considering the $D\rightarrow 4$ limit~\cite{Glavan:2019inb}. The similar idea was also considered  before motivated by the study of quantum corrections~\cite{Tomozawa:2011gp, Cognola:2013fva}. Many new predictions have been studied in the literature~\cite{Glavan:2019inb, Konoplya:2020bxa, Guo:2020zmf,  Fernandes:2020rpa, Casalino:2020kbt, Wei:2020ght, Konoplya:2020qqh, Kumar:2020owy, Hegde:2020xlv, Doneva:2020ped, Ghosh:2020vpc, Zhang:2020qew, Lu:2020iav, Singh:2020xju, Konoplya:2020ibi, Ghosh:2020syx, Konoplya:2020juj,  Kobayashi:2020wqy, Zhang:2020qam, HosseiniMansoori:2020yfj, Kumar:2020uyz, Singh:2020nwo, Wei:2020poh, Churilova:2020aca, Kumar:2020xvu, Islam:2020xmy,Liu:2020vkh, Heydari-Fard:2020sib, Konoplya:2020cbv, Jin:2020emq, Ai:2020peo, Zhang:2020sjh,EslamPanah:2020hoj, NaveenaKumara:2020rmi}. While the exact theory of the so called $4 D$ GB theory remains controversial~\cite{Gurses:2020ofy}, one concrete approach is to apply the Kaluza-Klein procedure, keeping consistently only the four-dimensional metric, together with the breathing mode.  One can then consistently set the dimensions $(D-4)$ of the internal space to zero,  leading to a scalar-tensor theory of the Hordenski class~\cite{Lu:2020iav}. This means that the procedure of the $D\rightarrow 4$ limit is not a simple trick.
So, a question arises naturally as to whether the emergent universe scenario can be realized in the  GB gravity by rescaling the GB coupling constant and considering the $D\rightarrow 4$ limit, against both homogeneous and inhomogeneous scalar perturbations simultaneously. We will answer this question in this paper.

 The remaining part of this paper is organized as follows. In Sec.~\ref{sec2}, we obtain the Einstein static universe  in the context of  the $4 D$ GB gravity with a perfect fluid. In Sec.~\ref{sec3}, we study the stability against both homogeneous and inhomogeneous scalar perturbations for spatially closed and open universes respectively.  Finally, we conclude  in Sec.~\ref{sec4}.


\section{Einstein static universe in $4D$ GB gravity}\label{sec2}

In this section, we will firstly briefly review the Friedmann equations in the GB gravity by rescaling the GB coupling constant and considering the $D\rightarrow 4$ limit. Then, we will give the Einstein static universe solution.
The action ${\cal S}$ of the $D$-dimensional GB gravity is given by
\be
{\cal S} =  \int d^D x \sqrt{-g} ( R + \alpha {\cal G } )  +{\cal S}_m \,,
\ee
where 
\be
{\cal G}= R^2 -4 R_{\mu\nu}R^{\mu\nu} +R_{\mu\nu\rho\sigma} R^{\mu\nu\rho\sigma} \,,
\ee
is the GB term, $\alpha$ is the coefficient of coupling constant and is positive in the heterotic string theory,  and $g$ and ${\cal S}_m$ respectively represent  the determinant of $g_{\mu\nu}$ and the action of matters. By varying $g^{\mu\nu}$, the EOM is given by
\be
G_{\mu\nu} + \alpha (-\frac12 g_{\mu\nu} {\cal G} +2 R_{\mu\lambda\rho\sigma} {R_\nu}^{\lambda\rho\sigma} -4 R_{\mu\lambda} {R_\nu}^\lambda -4 R_{\mu\rho\nu\sigma} R^{\rho\sigma} +2 R R_{\mu\nu}) =  T_{\mu\nu} \,, \label{eom1}
\ee
where $G_{\mu\nu}$ is the Einstein tensor, and  $T_{\mu\nu} = -\frac{1}{\sqrt{-g}}\frac{\partial {\cal S}_m}{\partial g^{\mu\nu}}$ is the energy-momentum tensor. The trace of the above equation is
\be
\frac{D-2}{2} R + \frac{(D-4) \alpha}{2} {\cal G} = {T^\mu}_\mu \,. \label{eom2}
\ee
First, we consider the $D$-dimensional FRW metric
\be
ds^2 = -dt^2 +a(t)^2 d\Omega_{D-1, \kappa}^2 \,, \label{nfrw}
\ee
where $a $ is the scale factor, $d\Omega_{D-1, \kappa}^2$ is the line element of $(D-1)$-dimensional maximally symmetric space  which can be written as
\be
d\Omega_{D-1, \kappa}^2 =\eta_{ij} dx^i dx^j =  \frac{dr^2}{1-\kappa r^2} +r^2 d\sigma^2 \,,
\ee
and $\kappa = 1, 0, -1$ denote the metric for the closed, flat and open universe respectively.  We assume that the matter is a perfect fluid, and the corresponding  energy-momentum tensor is given by
\be
T^{\mu\nu} = (\rho + P) u^\mu u^\nu +P g^{\mu\nu} \,, \quad \textup{with} \quad P = w \rho \,, \label{fluid}
\ee
where $\rho$ and $P$ represent energy density and pressure respectively, $w$ is the constant equation-of-state parameter, and velocity $u^\mu$ reads
\be
u^\mu = \left(1, 0, \dots, 0\right) \,, \quad \textup{satisfying} \quad u^\mu u_\mu = -1 \,. \label{velocity}
\ee
Substituting Eqs.(\ref{nfrw}) and (\ref{fluid}) into EOMs (\ref{eom1}) and (\ref{eom2}), we have the $D$-dimensional Friedmann equations 
 \bea
&& -\frac{(D-1)(D-2) (\kappa +\dot{a}^2 )}{2 a^2} - \frac{(D-1)(D-2)(D-3)(D-4)\alpha (\kappa +\dot{a}^2)^2}{2 a^4} = -\rho \,,\\
&&  -\frac{(D-1)(D-2)(D-3)(D-4)\alpha ((D-4)(\kappa +\dot{a}^2) +4 a \ddot{a})}{2 a^4} \nn \\
&&-\frac{(D-1)(D-2)((D-2)(\kappa+\dot{a}^2 ) +2 a\ddot{a} )}{2 a^2} = -\rho +(D-1) P \,.
 \eea
It is easy to see that the Gauss-Bonnet term vanishes in four dimensions. Now following  Ref.~\cite{Glavan:2019inb}, we consider 
\be
\alpha = \frac{\tilde{\alpha}}{D-4} \,, \label{rescale}
\ee
and then we adopt $D=4$. The metric reduces to the standard FRW metric, and the corresponding four-dimensional Friedmann equations become
\bea
&& \frac{3 \left(\dot{a}^2+\kappa \right)}{a^2} +\frac{3 \tilde{\alpha}  \left(\dot{a}^2+\kappa \right)^2}{a^4} =\rho \,,\\
&& -\frac{3 \left(2 a \ddot{a}+2 \left(\dot{a}^2+\kappa \right)\right)}{a^2} -\frac{12 \tilde{\alpha}  \ddot{a} \left(\dot{a}^2+\kappa \right)}{a^3}= -\rho +3 P \,,
\eea
where the dot represents the derivative with respect to time.
For the purpose of obtaining the Einstein static universe,  we let the scale factor $a(t)=a_0 = \textup{const}.$, and then the FRW metric reduces to the metric of the Einstein static universe
\be 
ds^2 = -dt^2 +a_0^2 \left[\frac{dr^2}{1-\kappa r^2} +r^2 \left(d\theta^2 +\sin^2\theta d\phi^2\right)\right] \,. \label{esu}
\ee
The solution is given by
\be
a_0 = \sqrt{\frac{{\tilde \alpha } \kappa (1-3 w)}{3 w+1}} \,,\qquad \rho = \frac{6 (3 w+1)}{{\tilde \alpha } (1-3 w)^2} \,.
\ee
First, it is easy to see that the solution does not exist in pure GR. Second, there is no Einstein static flat universe ($\kappa = 0$) in the four-dimensional GB theory filled with a perfect fluid. So we only discuss the  closed ($\kappa =1$) and open ($\kappa = -1$) Einstein static universes.


\section{Stability analysis}\label{sec3}


Now we study the stability of the Einstein static universe in the $4 D$ GB gravity. First we will study the linear EOMs in $D$ dimensions. Then we perform the $D\rightarrow 4$ limit. We use the symbols bar and tilde to represent background and perturbation components respectively.  The perturbed metric can be written as
\be
g_{\mu\nu} = \bar{g}_{\mu\nu} + h_{\mu\nu} \,,
\ee
where $\bar{g}_{\mu\nu}$ is the background metric which is given by~Eq.~(\ref{nfrw}) with $a = a_0$ and the $h_{\mu\nu}$ is a small perturbation.
For our purpose, we consider scalar perturbations in the Newtonian gauge in which ${h_\mu}^\nu$ is given by
\be
{h_\mu}^\nu = \textup{diag} \left(-2 \Psi \,, 2 \Phi \,, \dots \,, 2 \Phi\right) \,. \label{pert}
\ee
Now the indexes are lowered and raised by the background metric. By using the relation $g^{\mu\nu} g_{\nu\lambda} ={\delta^\mu}_\lambda$, the inverse metric is perturbed by
\be
\widetilde{g}^{\mu\nu} = -\bar{g}^{\mu\rho} \bar{g}^{\nu\sigma} h_{\rho\sigma} \,.
\ee
For a perfect fluid, the perturbations of energy density and pressure are $\widetilde{\rho}$ and $\widetilde{P} = w \widetilde{\rho}$ . The perturbations of velocity are given by
\be
  \widetilde{u}_0 = \widetilde{u}^0 = \frac{h_{00}}{2}   \,, \quad \widetilde{u}^i = \bar{g}^{ij} \, \widetilde{u}_j = \bar{g}^{ij} \bar{\nabla}_j U  \,,
\ee
where ``\,0\," and ``\,$i, j$\," denote time and space components respectively. The perturbed energy momentum tensor is
\be
\widetilde{T}^{\mu\nu} = P \, \widetilde{g}^{\mu\nu} +\widetilde{P} \, \bar{g}^{\mu\nu} +  (\widetilde{\rho} + \widetilde{P}) u^\mu u^\nu +(\rho + P) \widetilde{u}^\mu u^\nu +(\rho + P) u^\mu \widetilde{u}^\nu  \,,
\ee
where $u^\mu , \rho , and P$ represent the unperturbed components. Considering above expressions, the linearized equations of Eqs.~(\ref{eom1}) and (\ref{eom2}) become
\bea
&&\widetilde{G}_{\mu\nu} + \alpha (-\frac12 h_{\mu\nu} {\cal G} -\frac12 g_{\mu\nu} {\cal \widetilde{G}} +2 \widetilde{R}_{\mu\lambda\rho\sigma} {R_\nu}^{\lambda\rho\sigma} +2 R_{\mu\lambda\rho\sigma} {\widetilde{R}_\nu}^{\lambda\rho\sigma} -4 \widetilde{R}_{\mu\lambda} {R_\nu}^\lambda -4 R_{\mu\lambda} {\widetilde{R}_\nu}^\lambda \nn \\
&& -4 \widetilde{R}_{\mu\rho\nu\sigma} R^{\rho\sigma}  -4 R_{\mu\rho\nu\sigma} \widetilde{R}^{\rho\sigma} +2 \widetilde{R} R_{\mu\nu}+2 R \widetilde{R}_{\mu\nu}) =  \widetilde{T}_{\mu\nu} \,, \label{lineom1}\\
&& \frac{D-2}{2} \widetilde{R} + \frac{(D-4) \alpha}{2} {\cal \widetilde{G}} = {\widetilde{T}^\mu}_\mu \,.  \label{lineom2}
\eea
$\Psi, \Phi, \widetilde{\rho}$ and $U$ are functions of $(t, r, x_1, \dots, x_{D-2})$. For scalar perturbations, it is useful to perform a harmonic decomposition~\cite{Harrison:1967zza},
\bea
\Psi = \Psi_n(t)\, Y_n (r, x_1,\dots, x_{D-2}) \,,\quad \Phi = \Phi_n(t) \,Y_n (r, x_1,\dots, x_{D-2}) \,,\nn \\
\widetilde{\rho} = \rho \, \xi_n(t) \,Y_n (r, x_1,\dots, x_{D-2}) \,,\quad U = U_n(t) \,Y_n (r, x_1,\dots, x_{D-2}) \,.\label{decomp}
\eea
In these expressions, summations over co-moving wavenumber $n$  are implied. The harmonic function $Y_n (r, x_1,\dots, x_{D-2})$ satisfies~\cite{Harrison:1967zza}
\bea
\Delta Y_n (r, x_1,\dots, x_{D-2}) &=& \frac{1}{\sqrt{-\det{\eta}}}\frac{\partial}{\partial \xi^i} \left(\sqrt{-\det{\eta}} \eta^{ij} \frac{\partial  }{\partial \xi^j} Y_n (r, x_1,\dots, x_{D-2})\right) \nn \\
&=&  -k^2 Y_n (r, x_1,\dots, x_{D-2}) \,,
\eea
where $\Delta$ is the Laplacian operator, $\xi^i = (r, x_1,\dots, x_{D-2})$, and $k$ is a separation constant. For a spatially closed universe corresponding to $\kappa =1$, we have $k^2 =n(n+2)$ where the modes are discrete $(n= 0, 1, 2\dots)$. For a spatially open universe corresponding to $\kappa =-1$, we have  $k^2 =n^2+1$ where $n\ge 0$. Formally, $n = 0$ gives a spatially homogeneous mode and $n = 1, 2\cdots$ correspond to spatially inhomogeneous modes for both closed and open universes~\cite{Barrow:2003ni, Huang:2015kca}.
Substituting Eqs.~(\ref{pert}) and (\ref{decomp}) into (\ref{lineom1})-(\ref{lineom2}), after some algebra, we have the $D$-dimensional effective perturbed equations,
\bea
&&\frac{2 {\alpha}  (D-4) (D-3) (D-2) \kappa  \left(-a_0^2 (D-1) \ddot{\Phi}_n -(D-4) \Phi_n  \left(-D \kappa +\kappa +k^2\right)+k^2 \Psi_n \right)}{a_0^4} + \nn \\
&& \frac{(D-2) \left(k^2 \Psi_n -a_0^2 (D-1) \ddot{\Phi}_n -(D-2) \Phi_n  \left((1-D) \kappa +k^2\right) \right)}{a_0^2} =\rho  \xi_n ((D-1)  w -1) \,, \\ 
&&-\frac{\alpha  (D-4) (D-3) (D-2) \kappa  \left((D-1) \kappa  \Psi_n -2 \Phi_n  \left(-D \kappa +\kappa +k^2\right)\right)}{a_0^4} \nn\\ 
&&+\frac{(D-2) \left(\Phi_n \left(-D \kappa +\kappa +k^2\right)+(\kappa -D \kappa ) \Psi_n \right)}{a_0^2}+\rho  (2 \Psi_n -\xi_n )=0 \,,\\
&&a_0^2 \rho  (w+1) U_n -(D-2) \dot{\Phi}_n \left(a_0^2+2 \alpha  (D-4) (D-3) \kappa \right)  =0 \,, \\
&&\Psi_n  \left(a_0^2+2 \alpha  (D-4) (D-3) \kappa \right)-(D-3) \Phi_n  \left(a_0^2+2 \alpha  (D-4) (D-5) \kappa \right) =0 \,.
\eea
Now we rescale $\alpha$ as Eq.~(\ref{rescale}) and adopt $D= 4$, and then the corresponding perturbed equations are given by 
\bea
&&\frac{2 k^2 \Psi  \left(2 \tilde{\alpha}  \kappa +a_0^2\right)}{a_0^4}-\frac{4 \left(k^2-3 \kappa \right) \Phi }{a_0^2}+\left(-\frac{12 \alpha  \kappa }{a_0^2}-6\right) \ddot{\Phi} = \xi \rho  (3 w -1 ) \,, \\ 
&&-\frac{2 \tilde{\alpha}  \kappa  \left(3 \kappa  \Psi -2 \left(k^2-3 \kappa \right) \Phi \right)}{a_0^4}+\frac{2 \left(\left(k^2-3 \kappa \right) \Phi -3 \kappa  \Psi \right)}{a_0^2}+\rho  (2 \Psi -\xi ) =0 \,,\\
&&a_0^2 \rho  (w+1) U -2 \left(2 \tilde{\alpha } \kappa +a_0^2\right) \dot{\Phi } =0 \,, \\
&&\Psi \left(2 \tilde{\alpha}  \kappa +a_0^2\right)-\Phi  \left(a_0^2-2 \alpha  \kappa \right) =0 \,.
\eea
We can solve the equations to get
\bea
&&U_n =\frac{2 \left(2 \tilde{ \alpha}  \kappa +a_0^2\right) \dot{\Phi}_n }{a_0^2 \rho  (w+1)} \,, \\
&&\xi_n = -\frac{2 \Phi_n  \left(a_0^2 \left(k^2-6 \kappa \right)+2 \tilde{ \alpha}  \kappa  k^2\right)+6 a_0^2 \left(2 \tilde{ \alpha}  \kappa +a_0^2\right) \ddot{\Phi }_n }{a_0^4 \rho  (3 w-1)} \,, \\
&&\Psi_n = \frac{\Phi_n  \left(a_0^2-2 \tilde{ \alpha}  \kappa \right)}{2 \tilde{ \alpha}  \kappa +a_0^2} \,,
\eea
where $\Phi_n$  satisfies a second order ordinary differential equation
\be
\ddot{\Phi}_n +  Z \Phi_n =0 \,, \label{ode}
\ee
in which the dot represents the derivative with respect to time and 
\be
Z = -\frac{(3 w+1) \left(\kappa +3 k^2 w (w+1)-9 \kappa  w^2\right)}{3 \tilde{ \alpha}  \kappa  (w+1) (3 w-1)} \,.
\ee
To analyze the stability of the Einstein static universe in the GB gravity, we need to discuss the condition of existence of the oscillating solution of Eq.~(\ref{ode}). The condition is
\be
Z>0  \,.
\ee
Apart from this condition, we should require the coupling constant $\tilde{\alpha}$, the energy density of the perfect fluid $\rho$ and $a_0$ satisfy
\be
\tilde{\alpha} >0 \,, \quad \rho>0   \quad \textup{and} \quad a_0 > 0\,. 
\ee

In the case of $\kappa =1$, which corresponds to the Einstein static closed universe, the stability requires 
\bea
&& k^2 = 0 \,, \quad -\frac13 \le w \le \frac13 \,,\\
&& k^2\ge 8 \,, \quad f(k^2)<w<\frac{1}{3} \,,\quad \textup{with} \quad f(k^2) = \frac{\sqrt{9 k^4-12 k^2+36}-3 k^2}{6 \left(k^2-3\right)} \,.
\eea
Note that $f(k^2)$ is a monotonic increasing function, and as  $k^2 \rightarrow \infty$,  $f(k^2) \rightarrow 0$. As a result,  we have 
\be
-\frac13 < f(k^2\ge 8) < 0  \,.
\ee
So, the Einstein static closed universe can be stable against both the homogeneous ($k^2 = 0$) and inhomogeneous ($k^2 \ge 8$) scalar perturbations simultaneously under the condition
\be
 0 \le w<\frac{1}{3} \,.
 \ee
We find that the stable Einstein static closed universe filled with non-relativistic matter and  radiation does exist.

In the case of $\kappa = -1$, which corresponds to the Einstein static open universe, the stability requires 
\be
 w>\frac13 \,.
\ee

\section{Conclusion and discussion}\label{sec4}

We investigate the realization of the emergent universe scenario in the GB gravity by rescaling the GB coupling constant and considering the $D \rightarrow 4$ limit. 
First, we obtain the Einstein static universe filled with a perfect fluid in the $4 D$ GB gravity. Then, we find the stability conditions against both homogeneous and inhomogeneous scalar perturbations simultaneously  for spatially closed and open universes respectively. We find that the Einstein static universe with  closed spatial geometry can be stable against both the homogeneous and inhomogeneous scalar perturbations simultaneously. Such a stable universe is filled with ordinary matter and radiation. The Einstein static universe with  open spatial geometry can be also stable.

Finally, we want to point out that our work only shows that  the emergent universe scenario may be realized in the GB gravity by rescaling the GB coupling constant and considering the $D \rightarrow 4$ limit since the Einstein static universe is stable against scalar perturbations. The stabilities against vector and tensor perturbations are also worthy of investigation. Besides, it is also worth studying the stabilities in the concrete scalar tensor theory given in Ref.~\cite{Lu:2020iav}. Finally, there are also many other important issues in cosmology which should be studied in the novel four-dimensional theory. We leave these to the future work.


\section*{ACKNOWLEDGEMENTS}
We are grateful to Profs.~H. L\"u and Hao Wei for kind and useful discussions.  This work was supported in part by  NSFC grants No. 11435006, No. 11690034, No. 11775077, No. 11947216 and China Postdoctoral Science Foundation 2019M662785.

\renewcommand{\baselinestretch}{1.0}


\end{document}